OpenStudy (anonymous):

is there anyone that is good with matrices because i need help with this one

OpenStudy (anonymous):

Find the image matrix for the following counterclockwise rotation.

OpenStudy (anonymous):

Lets start with the transformation matrix for an arbitrary rotation: \[R =\left[\begin{array}{cc} \cos \theta & -\sin \theta \\ \sin \theta & \cos \theta \end{array}\right]\]

OpenStudy (anonymous):

A counterclockwise rotation of 180 degrees will simply be \[\theta = \pi\]

OpenStudy (anonymous):

ok so whats next

OpenStudy (anonymous):

\[R =\left[\begin{array}{cc} \cos \pi & -\sin \pi \\ \sin \pi & \cos \pi \end{array}\right] =\left[\begin{array}{cc} -1 & 0 \\ 0 & -1 \end{array}\right] \]

OpenStudy (anonymous):

Now apply the transformation to the coordinate vectors

OpenStudy (anonymous):

\[ \left[\begin{array}{cc} -1 & 0 \\ 0 & -1 \end{array}\right] \left[\begin{array}{cc} 3 \\ 1 \end{array}\right] = \left[\begin{array}{cc} -3 \\ -1 \end{array}\right] \]\[ \left[\begin{array}{cc} -1 & 0 \\ 0 & -1 \end{array}\right] \left[\begin{array}{cc} 6 \\ 2 \end{array}\right] = \left[\begin{array}{cc} -6 \\ -2 \end{array}\right] \]\[ \left[\begin{array}{cc} -1 & 0 \\ 0 & -1 \end{array}\right] \left[\begin{array}{cc} 4 \\ 5 \end{array}\right] = \left[\begin{array}{cc} -4 \\ -5 \end{array}\right] \] So we have the "image matrix" \[ \left[\begin{array}{cc} -3 & -6 & -4 \\ -1 & -2 & -5 \end{array}\right] \]

OpenStudy (anonymous):

ooooo ok i seee thank you so much